OpenStudy (anonymous):

Calculate the de Broglie wavelength in picometers (1 picometer =10^−12 meters) of an electron that has kinetic energy of 10 keV. The mass of an electron is 9.11 x 10^−31 kg

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@oOKawaiiOo

OpenStudy (anonymous):

The wavelength is given by \[ \lambda = \frac{h}{p}\] Using the kinetic energy with the relativistic energy equations you get \[K=E-mc^2\] where \[E = \gamma mc^2\] and \[\gamma = \frac{1}{\sqrt{1-\frac{v^2}{c^2}}}\] giving \[K=\gamma mc^2 - mc^2 \\ \quad =mc^2(\gamma -1) \\ \ \\ \quad K= mc^2 \left( \frac{1}{\sqrt{1-\frac{v^2}{c^2}}}-1 \right)\] which lets you solve for the velocity, v of the electron since you're given kinetic energy and rest mass of the electron. If you keep the velocity as a fraction of c until the final calculation, your life will be much easier! Then, again, our original equation for the wavelength is \[\lambda = \frac{h}{p}\] where relativistic momentum is \[p=\gamma mv\] leaving you with \[\lambda = \frac{h}{\gamma mv} \\ \ \\ \quad =\frac{h}{mv}\sqrt{1-\frac{v^2}{c^2}}\]

OpenStudy (anonymous):

A useful conversion for either your mass or your kinetic energy is \[1eV = 1.602 \times 10^{-19}J\] \[1eV/c^2 = 1.783 \times 10^{-36} kg\]

OpenStudy (anonymous):

Well, that all is excessive because it's not actually going that fast. It's probably sufficient to use \[K=\frac{p^2}{2m}\] to solve for p, then \[ \lambda = \frac{h}{p}\]

OpenStudy (anonymous):

Very nice. I was about to add a note that the velocity should be estimated first because it might be much less than c and no relativistic analysis would be needed.